OpenStudy (anonymous):

Divide the following polynomial using synthetic division: (x^4 - 3x^3 + 3x^2 - 3x + 6)/ (x - 2) @Hero Give me a hand? lol

OpenStudy (anonymous):

1 -3 2 -3 6 2 -2 0 -6 1 -1 0 -3 0 x 3 - x 2 - 3 = 0

OpenStudy (anonymous):

i give up on this one lol

OpenStudy (therealmeeeee):

ok i can make this easier for u to understand ok

OpenStudy (therealmeeeee):

it is easy first take x^5-8x^3 term

OpenStudy (therealmeeeee):

x^5 -8x^3 +24x^2 +12x +40 / (x+4) = x^4 +4x^3 x^5 +4x^4 ----------- 0 4x^4 -8x^3 4x^4 +16x^3 -------------- 0 -24x^3 so can you continue it ?

OpenStudy (therealmeeeee):

is this making sense

OpenStudy (anonymous):

yeah i can get it now makes more sense to me now. :)

OpenStudy (therealmeeeee):

lol